# ReinFlow: Fine-tuning Flow Matching Policy with Online Reinforcement Learning

### v6

@@ Abstract
We propose ReinFlow, a simple yet effective online reinforcement learning (RL) framework that fine-tunes a family of flow matching policies for continuous robotic control. Derived from rigorous RL theory, ReinFlow injects learnable noise into a flow policy’s deterministic path, converting the flow into a discrete-time Markov Process for exact and straightforward likelihood computation. This conversion facilitates exploration and ensures training stability, enabling ReinFlow to fine-tune diverse flow model variants stably, including Rectified Flow reflow and Shortcut Models shortcut , particularly at very few or even one denoising step. We benchmark ReinFlow in representative locomotion and manipulation tasks, including long-horizon planning with visual input and sparse reward. The episode reward of Rectified Flow policies obtained an average net growth of 135.36% after fine-tuning in challenging legged locomotion tasks while saving denoising steps and 82.63% of wall time compared to state-of-the-art diffusion RL fine-tuning method DPPO ren2024diffusionpolicypolicyoptimization . The success rate of the Shortcut Model policies in state and visual manipulation tasks achieved an average net increase of 40.34% after fine-tuning with ReinFlow at four or even one denoising step, whose performance is comparable to fine-tuned DDIM policies while saving computation time for an average of 23.20% . Code, model, and checkpoints available on the project website: https://reinflow.github.io/
@@ 1 Introduction
Recent years have seen rapid progress in the training of robotic models via imitation learning. Flow matching models, which combine the trifecta of precise modeling, fast inference, and minimal implementation, have emerged as a robust alternative to diffusion policies and a popular choice for robot action generation riemannian_flow_policy ; black2024pi0visionlanguageactionflowmodel ; zhang2025affordancebasedrobotmanipulationflow .
However, due to the scarcity of robot data and the embodiment gap, flow policies and many other imitation learning policies are often trained from datasets with mixed quality open_x_embodiment_rt_x_2023 , incurring suboptimal success rates even after supervised fine-tuning black2024pi0visionlanguageactionflowmodel . Although collecting more data may alleviate this problem, recent work data_scaling_laws_robot suggests that scaling the quantity of data may not be a panacea because, in a given environment, the success rate quickly plateaus when we only increase the number of demonstrations. Worse still, as an imitation learning method, flow policies lack a built-in exploration mechanism, implying that robots trained on imperfect data could struggle to complete challenging tasks where the agent needs to surpass expert demonstrations data_quality_IL .
Online reinforcement learning (RL) offers a promising solution to these challenges. By learning through trial and error, RL promises to overcome limitations associated with imperfect expert data and even achieve superhuman performance. Although recent studies have shown that it is possible to fine-tune diffusion policies via RL li2024learning ; ren2024diffusionpolicypolicyoptimization , training flow models with online RL remains technically challenging.
First, the theory has established that the refinement of a stochastic policy is vital for continuous control problems simchowitz2025pitfallsimitationlearningactions . However, for conditional flows, whose sample path is governed by a neural ordinary differential equation (ODE), even the log probability-a key factor that measures this stochasticity-could be challenging to obtain and unstable to backpropagate through. This problem is exacerbated when we infer flow policies at very few denoising steps, where the discretization error becomes large at a low inference cost. Second, compared to offline RL methods, online RL fine-tuning requires the policy to balance exploration and exploitation, especially in sparse reward settings pmlr-v205-guo23a . However, how to design a principled exploration mechanism remains elusive for conditional flows with a deterministic path.
@@
  • Algorithm Design. We propose ReinFlow, the first online RL algorithm to stably fine-tune a family of flow matching policies, especially in very few or even in one denoising step. We train a noise injection network to convert flows to a discrete-time Markov process with Gaussian transition probabilities for exact and tractable likelihood computation. Our design allows the noise net to automatically balance exploration with exploitation; it enjoys lightweight implementation, built-in exploration, and broad applicability to various variants of flow policy, including those parameterized with Rectified Flow reflow and Shortcut Models shortcut .
  • Empirical Validation. We perform extensive experiments in representative robot locomotion and manipulation tasks, with the agent receiving state or pixel observations and possibly accepting sparse rewards. Without reward shaping or scaling off-line data, our method, on average, improves the success rate over the pre-trained manipulation policy by 40.34% and increases the reward of locomotion policies by 135.36%. We achieve this improvement with a wall time reduction of 62.82% for all tasks compared to the state-of-the-art diffusion RL method DPPO ren2024diffusionpolicypolicyoptimization .
@@ 2 Related Work
Training diffusion-based policies from demonstrations has recently achieved impressive results in a variety of robot learning tasks ankile2024juicerdataefficientimitationlearning ; chi2024diffusionpolicyvisuomotorpolicy ; pearce2023imitatinghumanbehaviourdiffusion ; reuss2023goalconditionedimitationlearningusing ; sridhar2023nomadgoalmaskeddiffusion ; wang2024pocopolicycompositionheterogeneous ; ze20243ddiffusionpolicygeneralizable . However, their performance remains highly dependent on the quality of the demonstration data. In practice, demonstrations often contain mixed or suboptimal trajectories, which limits policy effectiveness and motivates the need for online fine-tuning.
Several offline diffusion-based RL methods can be extended to online settings. For example, Diffusion Q-Learning (DQL) wang2023diffusionpoliciesexpressivepolicy and Implicit Diffusion Q-Learning (IDQL) hansenestruch2023idqlimplicitqlearningactorcritic treat diffusion models as stochastic action policies and apply Q-learning updates. DIPO yang2023policyrepresentationdiffusionprobability adopts a similar idea, but employs a critic to directly refine the actions sampled using action gradients. Although these approaches leverage Q-function approximations to guide the diffusion actor, inaccurate Q estimates can introduce bias and destabilize training.
Alternatively, methods such as Q-Score Matching (QSM) psenka2025learningdiffusionmodelpolicy and Diffusion Policy Policy Optimization (DPPO) ren2024diffusionpolicypolicyoptimization fine-tune pre-trained diffusion policies using policy gradient techniques. Despite the growing interest in online RL for diffusion-based policies, by the time we release this work, few analogous methods exist for flow-based policies. The underlying mathematical distinctions between diffusion and flow models prevent direct adaptation of existing techniques liu2025flowgrpotrainingflowmatching , which poses additional challenges to develop effective online RL algorithms for flow policies.
@@
Flow matching models are more efficient than diffusion models, offering faster training, sampling, and better generalization lipman2023flow ; zheng2023guidedflowsgenerativemodeling . Following their success in imitation learning riemannian_flow_policy ; zhang2025affordancebasedrobotmanipulationflow and vision tasks lipman2023flow ; esser2024scaling ; kong2025hunyuanvideosystematicframeworklarge ; wan2025wanopenadvancedlargescale , recent work has begun integrating them into reinforcement learning (RL).
Flow Q-Learning (FQL) park2025flowqlearning trains flow policies offline and distills them for fine-tuning but lacks exploration during online adaptation, leading to suboptimal convergence. ReinFlow addresses this by injecting bounded, learnable noise into flow trajectories to promote exploration.
While Flow-GRPO liu2025flowgrpotrainingflowmatching and ORW-CFM-W2 fan2025onlinerewardweightedfinetuningflow also study online RL for flow models, they focus on vision tasks rather than continuous robotic control. ReinFlow instead provides a general policy gradient framework with a compact noise network and exact likelihood computation, unlike Flow-GRPO’s fixed noise or ORW-CFM-W2’s reward-weighted regression approach.
@@ 3 Problem Formulation
Flow matching  lipman2023flow transforms random variables from one distribution $p_{0}$ to another $p_{1}$ with flow mappings $\psi:[0,1]\times\mathcal{X}\to\mathcal{X}$, where $X_{t}:=\psi_{t}(X_{0}),t\in[0,1]$. This process is associated with an ODE: $\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}(X_{0})=v(t,\psi_{t}(X_{0}))$ where $X_{0}\sim p_{0}$. Rectified flow (ReFlow) reflow  [^1] For simplicity, we only consider 1-ReFlow and use 1-ReFlow and ReFlow in this work interchangeably. is a simple flow model with a straight ODE path $X_{t}=tX_{1}+(1-t)X_{0}$. The velocity field for Rectified Flow satisfies $v(t,X_{t})=\frac{\mathrm{d}}{\mathrm{d}t}X_{t}=X_{1}-X_{0}$, which implies that the training objective for Rectified Flow is given by
@@
Practitioners also sample $t$ from the beta distribution pizero or the logit normal distribution esser2024scaling . shortcut proposes “Shortcut Models” to further improve the generation quality of Rectified Flow at very few denoising steps by enforcing the velocity generated by two steps to align with that generated by a single step. During inference, we numerically solve the transport equation by integrating the learned velocity field: $\hat{X}_{1}=X_{0}+\sum_{k=0}^{K-1}v_{\theta}(t_{i},X_{t_{i}})\Delta t_{i}$, where $K$ is the number of denoising steps, $0=t_{0}<t_{1}<\ldots<t_{K-1}<1=t_{K}$ are the discretized time steps, $\Delta t_{i}=t_{i+1}-t_{i}$ is the step size.
@@
Flow policies offer significantly faster inference than autoregressive and diffusion models pizero ; zhang2025affordancebasedrobotmanipulationflow . A fast policy is valuable for robot learning, as a high inference frequency enhances the robot’s dexterity, and faster rollouts also reduce the wall-clock time of RL fine-tuning. Reducing denoising steps is arguably the most straightforward method to further accelerate flow policies, but when numerically solving the ODE, fewer integration steps increase the discretization error and reduce the quality of action generation sauer2011numerical . This work seeks to use RL to improve flow policies with minimal denoising steps, ideally a single step, to achieve the best of both worlds: rapid inference and high success rates.
@@ 4 Algorithm Design
2:Compute the advantage estimates $\widehat{A}_{i}:=\widehat{A}(o_{i},a_{i}^{K})$ by methods such as GAE schulman2018highdimensionalcontinuouscontrolusing
@@ 4.1 Likelihood Computation over a Short Denoising Trajectory
The log probability of a policy, which quantifies action stochasticity, is essential for policy gradient methods in continuous control sac ; ppo . However, computing the log probability for flow-matching policies with minimal denoising steps is challenging.
Although an exact log probability expression exists NEURIPS2018_69386f6b for continuous-time flow models,
@@
in practice, we need to simulate this integral with numerical solvers and estimate the divergence hutchinson1989stochastic :
@@
where $Z$ is a zero-mean random vector with an identity covariance matrix. However, the trace estimator involves Monte-Carlo error; Simulating the integral introduces discretization error, which becomes more pronounced with larger step sizes (fewer denoising steps) sauer2011numerical —a critical limitation for fast inference in robotic action generation. Treating the flow process at inference as a discrete-time Markov process can mitigate this issue. However, the intermediate variables follow a deterministic transition $p(X_{t+\Delta t}=x|X_{t})=\delta\left(x-X_{t}-v_{\theta}(t,X_{t})\Delta t\right)$, which renders the computation of the probability impossible.
@@ 4.2 Policy Gradient of a Discrete-time Markov Process
Theorem 4.1 builds on Theorem B.1 in qsm and established results in policy gradient theory agarwal2021theory . We provide the proofs in Appendix A.
Using Eq. (7), we enable the application of various modern deep policy gradient algorithms to optimize a policy parameterized by a discrete-time Markov process, including the noise-injected flows introduced in Section 4.1, which are of particular interest. Specifically, we can optimize the policy using either Eq. (9a) or Eq. (9b). In this work, we implement Eq. (9b) with the clipped surrogate loss ppo due to its stability.
@@
and employ off-policy methods, such as SAC sac , to update the policy.
In practice, we adopt “action chunking” during policy execution lai_huang_gershman_2022 , in which the policy outputs multiple actions in a batch given an observation. We explain how action chunking affects likelihood computation in Appendix A.2.
@@ 4.4 Regularization
One approach is to adopt Wasserstein regularization w2gan ; park2025flowqlearning , which constrains the Wasserstein-2 distance from the fine-tuned policy to the pre-trained policy. Studies show that $W_{2}$ regularization could improve training stability for generative models wgan ; w2gan . In practice, we minimize a tractable upper bound on this distance to enforce such constraint:
@@
In Eq. (12), $\Lambda_{o}$ indicates the set of distributions over $\mathcal{A}^{2}$ whose marginals are $\pi_{\theta_{\mathrm{old}}}(\cdot|o)$ and $\pi_{\theta}(\cdot|o)$. To implement Eq. (12), we replace the expectations with the sample mean over a batch of data. Inspired by park2025flowqlearning , we integrate $a$ and $a_{\mathrm{old}}$ from the same starting noise $a_{0}\sim\mathcal{N}(0,\mathbb{I}_{d_{A}})$, to control the stochasticity of the initial denoise action. We also note that when computing $\mathcal{R}_{W_{2}}$, we do not inject noise when sampling $a$ from $\pi_{\theta}$.
@@
Another technique is entropy regularization, which theory has shown accelerates convergence agarwal2019RLtheorybook ; global_convergence_NPG_entropy and encourages exploration sac for simple policy classes. For a flow matching policy parameterized with a discrete-time Markov process, we adopt the negative per-symbol entropy rate (or block entropy) schurmann1996entropy as the entropy regularizer. For a stationary POMDP and policy, we define the regularizer as
@@
Here, the last step is due to the definition of the noise-injected flow process in Eq. (6). $\mathbf{h}$ is the differential entropy operator cover1999elements_of_information_theory , of which normal distributions possess a simple closed-form expression specified in Section 3. According to schurmann1996entropy , the per-symbol entropy measures the Shannon entropy of a finite symbol sequence with long-range correlations. By minimizing $\mathcal{R}_{\mathbf{h}}$, we promote the agent to seek more diverse actions, enhancing exploration.

### v7

@@ Abstract
We propose ReinFlow, a simple yet effective online reinforcement learning (RL) framework that fine-tunes a family of flow matching policies for continuous robotic control. Derived from rigorous RL theory, ReinFlow injects learnable noise into a flow policy’s deterministic path, converting the flow into a discrete-time Markov Process for exact and straightforward likelihood computation. This conversion facilitates exploration and ensures training stability, enabling ReinFlow to fine-tune diverse flow model variants stably, including Rectified Flow Liu et al. (2022) and Shortcut Models Frans et al. (2024), particularly at very few or even one denoising step. We benchmark ReinFlow in representative locomotion and manipulation tasks, including long-horizon planning with visual input and sparse reward. The episode reward of Rectified Flow policies obtained an average net growth of 135.36% after fine-tuning in challenging legged locomotion tasks while saving denoising steps and 82.63% of wall time compared to state-of-the-art diffusion RL fine-tuning method DPPO Ren et al. (2024). The success rate of the Shortcut Model policies in state and visual manipulation tasks achieved an average net increase of 40.34% after fine-tuning with ReinFlow at four or even one denoising step, whose performance is comparable to fine-tuned DDIM policies while saving computation time for an average of 23.20% . Code, model, and checkpoints available on the project website: https://reinflow.github.io/
@@ 1 Introduction
Recent years have seen rapid progress in the training of robotic models via imitation learning. Flow matching models, which combine the trifecta of precise modeling, fast inference, and minimal implementation, have emerged as a robust alternative to diffusion policies and a popular choice for robot action generation Braun et al. (2024); Black et al. (2024a); Zhang and Gienger (2025).
However, due to the scarcity of robot data and the embodiment gap, flow policies and many other imitation learning policies are often trained from datasets with mixed quality Collaboration et al. (2023), incurring suboptimal success rates even after supervised fine-tuning Black et al. (2024a). Although collecting more data may alleviate this problem, recent work Lin et al. (2025) suggests that scaling the quantity of data may not be a panacea because, in a given environment, the success rate quickly plateaus when we only increase the number of demonstrations. Worse still, as an imitation learning method, flow policies lack a built-in exploration mechanism, implying that robots trained on imperfect data could struggle to complete challenging tasks where the agent needs to surpass expert demonstrations Belkhale et al. (2023).
Online reinforcement learning (RL) offers a promising solution to these challenges. By learning through trial and error, RL promises to overcome limitations associated with imperfect expert data and even achieve superhuman performance. Although recent studies have shown that it is possible to fine-tune diffusion policies via RL Li et al. (2024); Ren et al. (2024), training flow models with online RL remains technically challenging.
First, the theory has established that the refinement of a stochastic policy is vital for continuous control problems Simchowitz et al. (2025). However, for conditional flows, whose sample path is governed by a neural ordinary differential equation (ODE), even the log probability-a key factor that measures this stochasticity-could be challenging to obtain and unstable to backpropagate through. This problem is exacerbated when we infer flow policies at very few denoising steps, where the discretization error becomes large at a low inference cost. Second, compared to offline RL methods, online RL fine-tuning requires the policy to balance exploration and exploitation, especially in sparse reward settings Guo et al. (2023). However, how to design a principled exploration mechanism remains elusive for conditional flows with a deterministic path.
@@
  • Algorithm Design. We propose ReinFlow, the first online RL algorithm to stably fine-tune a family of flow matching policies, especially in very few or even in one denoising step. We train a noise injection network to convert flows to a discrete-time Markov process with Gaussian transition probabilities for exact and tractable likelihood computation. Our design allows the noise net to automatically balance exploration with exploitation; it enjoys lightweight implementation, built-in exploration, and broad applicability to various variants of flow policy, including those parameterized with Rectified Flow Liu et al. (2022) and Shortcut Models Frans et al. (2024).
  • Empirical Validation. We perform extensive experiments in representative robot locomotion and manipulation tasks, with the agent receiving state or pixel observations and possibly accepting sparse rewards. Without reward shaping or scaling off-line data, our method, on average, improves the success rate over the pre-trained manipulation policy by 40.34% and increases the reward of locomotion policies by 135.36%. We achieve this improvement with a wall time reduction of 62.82% for all tasks compared to the state-of-the-art diffusion RL method DPPO Ren et al. (2024).
@@ 2 Related Work
Training diffusion-based policies from demonstrations has recently achieved impressive results in a variety of robot learning tasks Ankile et al. (2024); Chi et al. (2024); Pearce et al. (2023); Reuss et al. (2023); Sridhar et al. (2023); Wang et al. (2024); Ze et al. (2024). However, their performance remains highly dependent on the quality of the demonstration data. In practice, demonstrations often contain mixed or suboptimal trajectories, which limits policy effectiveness and motivates the need for online fine-tuning.
Several offline diffusion-based RL methods can be extended to online settings. For example, Diffusion Q-Learning (DQL) Wang et al. (2023) and Implicit Diffusion Q-Learning (IDQL) Hansen-Estruch et al. (2023) treat diffusion models as stochastic action policies and apply Q-learning updates. DIPO Yang et al. (2023) adopts a similar idea, but employs a critic to directly refine the actions sampled using action gradients. Although these approaches leverage Q-function approximations to guide the diffusion actor, inaccurate Q estimates can introduce bias and destabilize training.
Alternatively, methods such as Q-Score Matching (QSM) Psenka et al. (2025) and Diffusion Policy Policy Optimization (DPPO) Ren et al. (2024) fine-tune pre-trained diffusion policies using policy gradient techniques. Despite the growing interest in online RL for diffusion-based policies, by the time we release this work, few analogous methods exist for flow-based policies. The underlying mathematical distinctions between diffusion and flow models prevent direct adaptation of existing techniques Liu et al. (2025), which poses additional challenges to develop effective online RL algorithms for flow policies.
@@
Flow matching models are more efficient than diffusion models, offering faster training, sampling, and better generalization Lipman et al. (2023); Zheng et al. (2023). Following their success in imitation learning Braun et al. (2024); Zhang and Gienger (2025) and vision tasks Lipman et al. (2023); Esser et al. (2024); Kong et al. (2025); Wan et al. (2025), recent work has begun integrating them into reinforcement learning (RL).
Flow Q-Learning (FQL) Park et al. (2025) trains flow policies offline and distills them for fine-tuning but lacks exploration during online adaptation, leading to suboptimal convergence. ReinFlow addresses this by injecting bounded, learnable noise into flow trajectories to promote exploration.
While Flow-GRPO Liu et al. (2025) and ORW-CFM-W2 Fan et al. (2025) also study online RL for flow models, they focus on vision tasks rather than continuous robotic control. ReinFlow instead provides a general policy gradient framework with a compact noise network and exact likelihood computation, unlike Flow-GRPO’s fixed noise or ORW-CFM-W2’s reward-weighted regression approach.
@@ 3 Problem Formulation
Flow matching  Lipman et al. (2023) transforms random variables from one distribution $p_{0}$ to another $p_{1}$ with flow mappings $\psi:[0,1]\times\mathcal{X}\to\mathcal{X}$, where $X_{t}:=\psi_{t}(X_{0}),t\in[0,1]$. This process is associated with an ODE: $\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}(X_{0})=v(t,\psi_{t}(X_{0}))$ where $X_{0}\sim p_{0}$. Rectified flow (ReFlow) Liu et al. (2022) [^1] For simplicity, we only consider 1-ReFlow and use 1-ReFlow and ReFlow in this work interchangeably. is a simple flow model with a straight ODE path $X_{t}=tX_{1}+(1-t)X_{0}$. The velocity field for Rectified Flow satisfies $v(t,X_{t})=\frac{\mathrm{d}}{\mathrm{d}t}X_{t}=X_{1}-X_{0}$, which implies that the training objective for Rectified Flow is given by
@@
Practitioners also sample $t$ from the beta distribution Black et al. (2024b) or the logit normal distribution Esser et al. (2024). Frans et al. (2024) proposes “Shortcut Models” to further improve the generation quality of Rectified Flow at very few denoising steps by enforcing the velocity generated by two steps to align with that generated by a single step. During inference, we numerically solve the transport equation by integrating the learned velocity field: $\hat{X}_{1}=X_{0}+\sum_{k=0}^{K-1}v_{\theta}(t_{i},X_{t_{i}})\Delta t_{i}$, where $K$ is the number of denoising steps, $0=t_{0}<t_{1}<\ldots<t_{K-1}<1=t_{K}$ are the discretized time steps, $\Delta t_{i}=t_{i+1}-t_{i}$ is the step size.
@@
Flow policies offer significantly faster inference than autoregressive and diffusion models Black et al. (2024b); Zhang and Gienger (2025). A fast policy is valuable for robot learning, as a high inference frequency enhances the robot’s dexterity, and faster rollouts also reduce the wall-clock time of RL fine-tuning. Reducing denoising steps is arguably the most straightforward method to further accelerate flow policies, but when numerically solving the ODE, fewer integration steps increase the discretization error and reduce the quality of action generation Sauer (2011). This work seeks to use RL to improve flow policies with minimal denoising steps, ideally a single step, to achieve the best of both worlds: rapid inference and high success rates.
@@ 4 Algorithm Design
2:Compute the advantage estimates $\widehat{A}_{i}:=\widehat{A}(o_{i},a_{i}^{K})$ by methods such as GAE Schulman et al. (2018)
@@ 4.1 Likelihood Computation over a Short Denoising Trajectory
The log probability of a policy, which quantifies action stochasticity, is essential for policy gradient methods in continuous control Haarnoja et al. (2018); Schulman et al. (2017). However, computing the log probability for flow-matching policies with minimal denoising steps is challenging.
Although an exact log probability expression exists Chen et al. (2018) for continuous-time flow models,
@@
in practice, we need to simulate this integral with numerical solvers and estimate the divergence Hutchinson (1989):
@@
where $Z$ is a zero-mean random vector with an identity covariance matrix. However, the trace estimator involves Monte-Carlo error; Simulating the integral introduces discretization error, which becomes more pronounced with larger step sizes (fewer denoising steps) Sauer (2011)—a critical limitation for fast inference in robotic action generation. Treating the flow process at inference as a discrete-time Markov process can mitigate this issue. However, the intermediate variables follow a deterministic transition $p(X_{t+\Delta t}=x|X_{t})=\delta\left(x-X_{t}-v_{\theta}(t,X_{t})\Delta t\right)$, which renders the computation of the probability impossible.
@@ 4.2 Policy Gradient of a Discrete-time Markov Process
Theorem 4.1 builds on Theorem B.1 in Psenka et al. (2024) and established results in policy gradient theory Agarwal et al. (2021). We provide the proofs in Appendix A.
Using Eq. (7), we enable the application of various modern deep policy gradient algorithms to optimize a policy parameterized by a discrete-time Markov process, including the noise-injected flows introduced in Section 4.1, which are of particular interest. Specifically, we can optimize the policy using either Eq. (9a) or Eq. (9b). In this work, we implement Eq. (9b) with the clipped surrogate loss Schulman et al. (2017) due to its stability.
@@
and employ off-policy methods, such as SAC Haarnoja et al. (2018), to update the policy.
In practice, we adopt “action chunking” during policy execution Lai et al. (2022), in which the policy outputs multiple actions in a batch given an observation. We explain how action chunking affects likelihood computation in Appendix A.2.
@@ 4.4 Regularization
One approach is to adopt Wasserstein regularization Korotin et al. (2021); Park et al. (2025), which constrains the Wasserstein-2 distance from the fine-tuned policy to the pre-trained policy. Studies show that $W_{2}$ regularization could improve training stability for generative models Arjovsky et al. (2017); Korotin et al. (2021). In practice, we minimize a tractable upper bound on this distance to enforce such constraint:
@@
In Eq. (12), $\Lambda_{o}$ indicates the set of distributions over $\mathcal{A}^{2}$ whose marginals are $\pi_{\theta_{\mathrm{old}}}(\cdot|o)$ and $\pi_{\theta}(\cdot|o)$. To implement Eq. (12), we replace the expectations with the sample mean over a batch of data. Inspired by Park et al. (2025), we integrate $a$ and $a_{\mathrm{old}}$ from the same starting noise $a_{0}\sim\mathcal{N}(0,\mathbb{I}_{d_{A}})$, to control the stochasticity of the initial denoise action. We also note that when computing $\mathcal{R}_{W_{2}}$, we do not inject noise when sampling $a$ from $\pi_{\theta}$.
@@
Another technique is entropy regularization, which theory has shown accelerates convergence Agarwal et al. (2019); Cen et al. (2022) and encourages exploration Haarnoja et al. (2018) for simple policy classes. For a flow matching policy parameterized with a discrete-time Markov process, we adopt the negative per-symbol entropy rate (or block entropy) Schürmann and Grassberger (1996) as the entropy regularizer. For a stationary POMDP and policy, we define the regularizer as
@@
Here, the last step is due to the definition of the noise-injected flow process in Eq. (6). $\mathbf{h}$ is the differential entropy operator Cover (1999), of which normal distributions possess a simple closed-form expression specified in Section 3. According to Schürmann and Grassberger (1996), the per-symbol entropy measures the Shannon entropy of a finite symbol sequence with long-range correlations. By minimizing $\mathcal{R}_{\mathbf{h}}$, we promote the agent to seek more diverse actions, enhancing exploration.